OpenStudy (anonymous):

During each workout, he jogged and he rode his bike. He rode his bike for 45 minutes each day. He worked out for a total of 360 minutes last week. What is the number of minutes that Allen jogged each day last week? Enter your answer in the box below.

OpenStudy (anonymous):

help pls

OpenStudy (dangerousjesse):

\[360=45+j \]\[j=360-45\]

OpenStudy (dangerousjesse):

Solve for j first.

OpenStudy (anonymous):

i got 315

OpenStudy (dangerousjesse):

Yes, now \[45*7=?\]

OpenStudy (anonymous):

315

OpenStudy (anonymous):

45x7=315

OpenStudy (dangerousjesse):

So what's 45/7?

OpenStudy (anonymous):

like a fraction

OpenStudy (anonymous):

or divide

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

It could be both; \[\frac{x}{y}=x \div y\]

OpenStudy (dangerousjesse):

But division would be easier

OpenStudy (anonymous):

do u know the answer

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

Yes, 27 minutes of jogging every day (if they mean a 5-day week). That's what's confusing me.

OpenStudy (anonymous):

oh ok thanks yo

OpenStudy (anonymous):

can u kik me sebastiandope pls man i need help for other questions

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (anonymous):

lol i dident mean yo as a man but can u help me on here

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

Sure, I can help you here.

OpenStudy (anonymous):

Marla bought one necklace and two bracelets for $67.50. Use the following information to determine the amount she paid for each bracelet: The price of the necklace was $27. The bracelets were both the same price.

OpenStudy (anonymous):

do u know the answer for that pls

OpenStudy (dangerousjesse):

\[27+2x=67.5 \]

OpenStudy (anonymous):

so how much did she pay for each braclet

OpenStudy (anonymous):

she paid 27 right ?

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

Solve for x: \[2 x+27 = 67.5 \]Isolate terms with x to the left hand side. Subtract 27 from both sides: \[2 x+(27-27) = 67.5-27 \]Look for two terms that sum to zero. \[27-27 = 0: \]\[2 x = 67.5-27 \]Evaluate \[67.5-27. \] \[67.5-27 = 40.5: \]\[2 x = 40.5 \]Divide both sides by a constant to simplify the equation. Divide both sides of \[2 x = 40.5~~ by~~ 2:\] (2 x)/2 = 40.5/2 Any nonzero number divided by itself is one. \[2\div2 = 1:\]\[x = 40.5\div2\]Express 40.5/2 in decimal form. \[40.5 \div 2 = 20.25: \]

OpenStudy (dangerousjesse):

Which leaves you with...?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

right ?

OpenStudy (dangerousjesse):

No, that's how much she paid for the necklace. We're trying to find the price of the bracelets.

OpenStudy (anonymous):

can u give me the answer pls

OpenStudy (dangerousjesse):

If \[67.50-27=40.5\]then what's\[40.5 \div 2?\]

OpenStudy (anonymous):

20.25

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

Yes, good job :)

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

Two equations are shown: Equation 1: 3 4 (x−12)=12 Equation 2: 3 4 y−12=12 Solve each equation. Then, enter a number in each box to make this statement true. The value of x is: The value of y is:

OpenStudy (anonymous):

i got 24 for both of them idk if im right

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (anonymous):

No it is 16